the point on the curve (x)^(1/2)+(y)^1/2)=2*(a)^(2) where the tangents are equally inclined to the axes,is


plz ansr this question briefly o that i can understand it easily.

Grade:

1 Answers

vikas askiitian expert
13 years ago

eq of curve is x^1/2  +  y^1/2 =a^2

 

  let (x, y,) be the point on the curve at which we have to draw  the tangent

 

 since (x, y,) lies on the curve so it will satisfy the eq of curve

   so x,^1/2 + y,^1/2 =a^2  ..................eq1 

   slope of tangent =dy/dx= -(y,/x,)^1/2 at (x, y,)

   tangent is equally inclined to axes therefore slope =1 or -1....eq3

   solving eq 1 ,2 and 3 

  required point comes out to be (a^4/4 , a^4/4)